# Gaussian Differentially Private Human Faces Under a Face Radial Curve Representation

### v1

@@ 1 Introduction
In statistical analyses, data and parameters appear in varying degrees of complexity, from simpler forms such as scalars and vectors to more complex such as spherical or hyperbolic, for instance. The structural constraints inherent to data need to be respected throughout any analysis as has been shown in the “intrinsic statistics" frameworks (Pennec, 2006; Bhattacharya and Patrangenaru, 2003) for accurate estimation and to preserve said structure. Complex data structures tend to come hand in hand with complex statistical computations and hence the techniques for handling such data have not been widely studied outside of specific scenarios. Further, the sheer amount of data that is captured from individuals has increased significantly, and of course, this produces a growing concern for one’s “privacy". In this paper, to support broader sharing of confidential data, we propose releasing a Gaussian Differentially Private, GDP, average 3D human face.
@@
Motivation and Related Literature: Data that live in nonlinear spaces can be challenging to work within the DP framework, as shown in the context of manifolds in Reimherr et al. (2021); Soto et al. (2022); Utpala et al. (2022), private Riemannian optimization in Han et al. (2022), and Gaussian DP on manifolds Jiang et al. (2023). Preservation of structure has also been considered in the context of private covariance estimation for linear regression in Sheffet (2019) and private principal component analysis in Chaudhuri et al. (2013) which is connected to the Stiefel manifold. For privacy, our proposed “face radial curve" representation of a human face are functions extracted from a disk and hence lends itself to be examined under the lens of private functional data analysis (FDA) which has been considered inWasserman and Zhou (2010) and Mirshani et al. (2017), and references therein.
@@
With our faces constantly being captured (e.g., at a grocery store self-checkout), one could question the privacy protections in place of the collected data. Further, there is a vast amount of literature on identification of individuals from facial data in the area of “face recognition," see, for instance, the expansive literature review in Kortli et al. (2020). Also, perhaps quite surprisingly, Venkatesaramani et al. (2021) and Klimentidis and Shriver (2009) showed that one could identify genomic information from a person’s 2D face image; the former study further showed that adding noise to said 2D images can help protect this re-identification. We do not work with 2D images here, however, the need for privacy is still present. The need for privacy for 3D faces is similar to that in 2D images. Anthropological studies are often interested in identification of DNA using labeled 3D faces (Sero et al., 2019), the connections between DNA genotype and the associated facial phenotype (White et al., 2021; Naqvi et al., 2021; Weinberg et al., 2019), and average faces for demographics such as age and race classification (Tokola et al., 2015). In such studies, one might want to release the data, or may even be required to, so offering provable confidentiality guarantees for people in the studies becomes an important task.
@@
To generate our representation we rely on tools from shape analysis. The earlier forms of shape analysis, such as Kendall’s shape space Kendall (1984), are limited to only representing a shape as a finite point cloud. The field has expanded since to consider more complex data structures such as continuous curves, both planar and space, as in Trouvé and Younes (2005); Klassen et al. (2004); Srivastava et al. (2010), and surfaces as in Jermyn et al. (2017); Su et al. (2020). Human faces have been considered in this space in a similar, yet subtly different, manner such as in Samir et al. (2006); Drira et al. (2010) in which faces are represented as a set of curves “facial curves" and “radial curves", respectively. These methods represent a face with curves that are generated independently of each other while we use an entire disk-parameterization to capture features across all faces simultaneously.
@@
Main Contributions: We develop a novel representation for a collection of 3D faces via a set of curves which we extract from disk parameterizations and refer to these as face radial curves. We construct the face radial curves using tools from statistical shape analysis in the interest of preserving the shape of an average face during the data sanitization process. Further, we extend existing approximate DP FDA tools into the Gaussian DP framework (Dong et al., 2019), a recent notion of privacy with attractive properties including “tight" composition. Under our FDA Gaussian DP mechanism, we generate a private average face of a collection of faces under our representation. While we use our mechanism and representation for faces to address the inherent data privacy concerns, this same methodology can be applied to any surfaces diffeomorphic to a unit disk such as terrain models and additionally any settings where one wants private functional statistical summaries.
@@ 2 Notation and Background on Face Representation
We require a disk-parameterized representation of each face from which we extract a set of functions. We parameterize each face independently but “align" and “register" each face to a template. Here, we broadly describe these relevant techniques from shape analysis including their necessity. To the best of our knowledge, the use of parameterized surfaces in the context of DP has not been explored. The software to accomplish parameterization, registration, and alignment are fully described in Jermyn et al. (2017), with accompanying implementation at GitHub repository (Laga, 2022).
@@ 2.1 Parameterization
We first compute a Delaunay triangulation using the MeshLab software (Cignoni et al., 2008) yielding a triangulated mesh. Given a point cloud $X\in\mathbb{R}^{p\times 3}$ the triangulated mesh, $M=\{V,E,T\}$, is an object which consists of a set of vertices $V$ (the original points), edges $E$, and triangles $T$ composed of said vertices and edges. The triangles meet at edges, do not overlap, and jointly represent a mesh or surface. The left panel of Figure 1 displays a face as a triangulated mesh. We generate a disk conformal map of $M$ as in Choi and Lui (2018); a disk conformal map of the triangulated mesh is a mapping from said mesh to a disk which preserves the angles of the triangles or, more generally, the local geometry of the mesh.
Lastly, we generate a disk-parameterized surface as in Jermyn et al. (2017); Laga et al. (2018) via the disk conformal mapping. Figure 1 displays two disk-parameterized surfaces in the middle and right panels; we expand on the differences between these two panels in A.6. Each face, $f$, is thus a map $f:D\rightarrow\mathbb{R}^{3}$ with $D=\{r,\theta|0\leq r\leq 1,0\leq\theta<2\pi\}$, the unit disk.
@@ 3 Novel Face Radial Curves Representation
We parameterize the entire face using a disk, however, our goal is to work with what we refer to as “face radial curves." This concept is similar to that of facial curves (Samir et al., 2006) and radial curves (Drira et al., 2010). In the facial recognition literature, for instance, to construct a radial curve representation one picks a focal point, typically the tip of the nose, and overlays curves on the surface of the face with the nose as the center and each point of the curve being equidistant from the tip of the nose. That is, given a focal point $p$ each radial curve is $g=\{x|r=d_{F}(p,x)\}$ for a given $r\geq 0$ and where $d_{F}$ is distance measured on the surface of the face.
These previously mentioned methodologies do not take into account that the distribution of features on faces is not uniform for everyone. That is, let $g_{ij}$ represent the $j$th radial curve of the $i$th face, then $g_{ij}$ may be the curve which goes directly on the middle of the eyes for individual $i$ but $g_{i^{\prime}j}$ may be directly on the forehead of individual $i^{\prime}$ for the same $j$. Features, such as eyes and the mouth, being disproportionately farther or closer to the focal point or each other hence causes issues. These methodologies do, however consider registration within each $j$. Therefore, registration has been considered within each set of curves $j$ but not across all $j$s simultaneously. We propose a new method to have an entire global alignment and registration across all sets of curves and faces.
@@
We apply our proposed method to data described in Sero et al. (2019); more details are available A.4. The left panel of Figure 2 displays a face as a point cloud and the right panel is the same face but with the disk parameterization and some of the face radial curves overlain. While the disk parameterization is an infinite collection of these curves, we pick some number of these curves to represent the face. The number of curves is a tuning parameter where more curves means more definition in the face but only to an extent. We discuss this further in 5.
@@ 4 Gaussian Differential Privacy for Functional Data
Since the conception of DP (Dwork et al., 2006), noise calibration has been considered from different perspectives leading to many variants. For instance, “zero concentrated differential privacy" (Bun and Steinke, 2016) and “Rényi differential privacy" (Mironov, 2017) are notions of DP from the perspective of a divergence of the distribution of the mechanism. In the present paper we utilized “Gaussian DP" (GDP, Dong et al. (2019)), which considers DP from the perspective of a particular set of hypothesis tests. As one of our key contributions in this paper, we extend GDP to functional data analysis (FDA) in Section 4.2. We begin with a brief overview of DP for FDA.
@@ 4.1 Background on Differential Privacy for Functional Data
In this section and in A.5, we describe approximate DP, $(\epsilon,\delta)$-DP, in the context of FDA. This is a needed background for our proposed extension of GDP into the space of functional data. For a more thorough exposition on DP, see Blum et al. (2005); Dwork and Roth (2014); Dwork et al. (2006).
@@
We consider releasing an estimate of a private mean function. To sanitize functions we rely on tools and foundations of functional data analysis in the domain of privacy as in Hall et al. (2013); Alda and Rubinstein (2017); Mirshani et al. (2017), with the latter considering spaces more extensive than functions. The infinite dimensional nature of function valued summaries presents a critical challenge in establishing formal privacy. In particular, traditional probability densities become much more complicated as there is no default or baseline measure in infinite dimensions (unlike Lebesgue measure in $\mathbb{R}^{d}$). To overcome this challenge, there are currently two approaches. The first, taken by Hall et al. (2013) is to work in finite dimensions and then take careful limits. The second approach, introduced in Mirshani et al. (2017) and which we follow here, is to utilize carefully constructed infinite dimensional densities so that the probability inequalities can be worked out directly (and avoid having to take limits).
@@
Now that we have well defined noise, we consider realeasing a specific summary statistic: a private mean function with sample mean $h(D)=\bar{X}=\frac{1}{n}\sum_{i}x_{i}$. To have some control on the smoothness of this estimate, we enforce smoothness by relying on an optimization problem with a penalty. That is, we let $h(D)=\operatorname{argmin}_{m\in\mathcal{H}}\sum_{i}\|x_{i}-m\|^{2}_{\mathbb{H}}+\phi\|m\|_{\mathcal{H}}^{2}$ with $\phi$ being a smoothness penalty parameter. This approach ensures the required noise compatibility. For our summary statistic, the sensitivity, $\sup_{D\sim D^{\prime}}\|h(D)-h(D^{\prime})\|^{2},$ can be shown to be bounded as $\Delta^{2}\leq 4\tau^{2}/(n^{2}\phi)$ (Mirshani et al., 2017), where $\tau$ is a finite bound on the norm in the $\mathbb{H}$ space of the elements of all $D\in\mathcal{D}$.
@@ 4.2 Extension of GDP to Functional Data
One especially useful interpretation of approximate DP comes from Wasserman and Zhou (2010), which relates DP to hypothesis testing. In particular, for a sanitized output, $\tilde{h}(D)$, one can consider statistical tests for determining if the true underlying data source was $D$ or some adjacent data set $D^{\prime}$. In these simple cases, the optimal test is well known (Neyman-Pearson Lemma), so one can talk about the optimal type 2 error (1-power) of a statistical test for given type 1 error rate. It turns out that DP gives a bound for the type 2 error relative to the a type 1 error, which Dong et al. (2019) took a step further to define Gaussian DP. In particular, Dong et al. (2019) show a mechanism is $\mu$-GDP if its entire type 1/type 2 error tradeoff curve is bounded by that of a curve coming from comparing a $N(0,1)$ to $N(\mu,0)$, meaning it is harder to distinguish 00 and $\mu$ from $N(0,1)$ than it is $h(D)$ from $h(D^{\prime})$ from $\tilde{h}(D)$. We provide the formal definition below.
@@
Corollary 4.2 also appears in Balle and Wang (2018) in the context of the calibrating the Gaussian mechanism, but in the presented form the corollary is more readily applicable for our setting.
@@
The last equality is a reparametrization setting $\mu=\Delta/\sigma$ and the rest of the details are in the A.7. ∎
For facial radial curves and other disc-like surfaces, by construction, we need to sanitize many mean curves, so we require composition of a multitude of mechanisms with their respective budgets $\mu_{i}$’s. Composition in the GDP framework is “tight." Given two mechanisms $\tilde{h}_{1}$, $\tilde{h}_{2}$ with privacy parameters $\mu_{1}$ and $\mu_{2}$, their composition $(\tilde{h}_{1},\tilde{h}_{2})$ is $\sqrt{\mu_{1}^{2}+\mu_{2}^{2}}$-GDP (Dong et al., 2019, Corallary 3.3).
@@ 5 Face Radial Curve Example
By design we have closed curves, so to retain this structure we use a periodic kernel that takes the form $k(s,t)=\exp\left(-\left[d_{\mathbb{S}^{1}}(\omega(t),\omega(s))/\rho\right]^{\alpha}\right)$ (Gneiting, 2013) with $t,s\in[0,1]$. The distance of two points $a,b$ on $\mathbb{S}^{1}$ is $\arccos\langle a,b\rangle$, however, the previous $t,s$ are parameters on the unit interval, so we first “wrap" the interval around the circle to compute this distance hence the need for $\omega(\cdot)$, a wrapping function. This kernel is a powered exponential on spheres with kernel range parameter $\rho$ and smoothness parameter $\alpha\in(0,1]$. For our experiments we set $\alpha=1$ as we can control smoothness instead via $\phi$ in the mean estimation as in 4.1.

### v2

@@ 1 Introduction
In statistical analyses, data and parameters appear in varying degrees of complexity, from simpler forms such as scalars and vectors to more complex such as spherical or hyperbolic, for instance. The structural constraints inherent to data need to be respected throughout any analysis as has been shown in the “intrinsic statistics” frameworks (Pennec, 2006; Bhattacharya & Patrangenaru, 2003) for accurate estimation and to preserve said structure. Complex data structures tend to come hand in hand with complex statistical computations and hence the techniques for handling such data have not been widely studied outside of specific scenarios. Further, the sheer amount of data that is captured from individuals has increased significantly, and of course, this produces a growing concern for one’s “privacy”. In this paper, to support broader sharing of confidential data, we propose releasing a Gaussian Differentially Private, GDP, average 3D human face.
@@
Motivation and Related Literature: Data that live in nonlinear spaces can be challenging to work within the DP framework, as shown in the context of manifolds in Reimherr et al. (2021); Soto et al. (2022); Utpala et al. (2022), private Riemannian optimization in Han et al. (2022), and Gaussian DP on manifolds Jiang et al. (2023). Preservation of structure has also been considered in the context of private covariance estimation for linear regression in Sheffet (2019) and private principal component analysis in Chaudhuri et al. (2013) which is connected to the Stiefel manifold. For privacy, our proposed “face radial curve” representation of a human face are functions extracted from a disk and hence lends itself to be examined under the lens of private functional data analysis (FDA) which has been considered inWasserman & Zhou (2010) and Mirshani et al. (2017), and references therein.
@@
With our faces constantly being captured (e.g., at a grocery store self-checkout), one could question the privacy protections in place of the collected data. Further, there is a vast amount of literature on identification of individuals from facial data in the area of “face recognition,” see, for instance, the expansive literature review in Kortli et al. (2020). Also, perhaps quite surprisingly, Venkatesaramani et al. (2021) and Klimentidis & Shriver (2009) showed that one could identify genomic information from a person’s 2D face image; the former study further showed that adding noise to said 2D images can help protect this re-identification. For 2D images, two common practices to attain privacy of faces are blurring and pixelation (Li & Choi, 2021; Vishwamitra et al., 2017). A major goal in these previous works is privacy in form of anonymity which is usually measured by re identification of individuals; our work, however, defines privacy by satisfying the conditions of differential privacy. This distinction is necessary as the former, generally, considers successful privacy by lack of re identification while preserving utility while for the latter differential privacy is a property of our mechanism. We do not work with 2D images here, however, the need for privacy is still present. The need for privacy for 3D faces is similar to that in 2D images. Anthropological studies are often interested in identification of DNA using labeled 3D faces (Sero et al., 2019), the connections between DNA genotype and the associated facial phenotype (White et al., 2021; Naqvi et al., 2021; Weinberg et al., 2019), and average faces for demographics such as age and race classification (Tokola et al., 2015). In such studies, one might want to release the data, or may even be required to, so offering provable confidentiality guarantees for people in the studies becomes an important task.
@@
To generate our representation we rely on tools from shape analysis. The earlier forms of shape analysis, such as Kendall’s shape space (Kendall, 1984), are limited to only representing a shape as a finite point cloud. The field has expanded since to consider more complex data structures such as continuous curves, both planar and space, as in Trouvé & Younes (2005); Klassen et al. (2004); Srivastava et al. (2010), and surfaces as in Jermyn et al. (2017); Su et al. (2020). Human faces have been considered in this space in a similar, yet subtly different, manner such as in Samir et al. (2006); Drira et al. (2010) in which faces are represented as a set of curves “facial curves” and “radial curves”, respectively. These methods represent a face with curves that are generated independently of each other while we use an entire disk-parameterization to capture features across all faces simultaneously.
@@
Main Contributions: We develop a novel representation for a collection of 3D faces via a set of curves which we extract from disk parameterizations and refer to these as face radial curves. We construct the face radial curves using tools from statistical shape analysis in the interest of preserving the shape of an average face during the data sanitization process. Further, we extend existing approximate DP FDA tools into the Gaussian DP framework (Dong et al., 2019), a recent notion of privacy with attractive properties including “tight” composition. Under our FDA Gaussian DP mechanism, we generate a private average face of a collection of faces under our representation. While we use our mechanism and representation for faces to address the inherent data privacy concerns, this same methodology can be applied to any surfaces diffeomorphic to a unit disk such as terrain models and additionally any settings where one wants private functional statistical summaries.
@@ 2 Notation and Background on Face Representation
We require a disk-parameterized representation of each face from which we extract a set of functions. We parameterize each face independently but “align” and “register” each face to a template. Here, we broadly describe these relevant techniques from shape analysis including their necessity. To the best of our knowledge, the use of parameterized surfaces in the context of DP has not been explored. The software to accomplish parameterization, registration, and alignment are fully described in (Jermyn et al., 2017), with accompanying implementation at GitHub repository (Laga, 2022). Further, in Figure 10 we display the entire pipeline for our methodology.
@@ 2.1 Parameterization
We first compute a Delaunay triangulation using the MeshLab software (Cignoni et al., 2008) yielding a triangulated mesh. Given a point cloud $X\in\mathbb{R}^{p\times 3}$ the triangulated mesh, $M=\{V,E,T\}$, is an object which consists of a set of vertices $V$ (the original points), edges $E$, and triangles $T$ composed of said vertices and edges. The triangles meet at edges, do not overlap, and jointly represent a mesh or surface. The left panel of Figure 1 displays a face as a triangulated mesh. We generate a disk conformal map of $M$ as in Choi & Lui (2018); a disk conformal map of the triangulated mesh is a mapping from said mesh to a disk which preserves the angles of the triangles or, more generally, the local geometry of the mesh.
Lastly, we generate a disk-parameterized surface as in Jermyn et al. (2017); Laga et al. (2018) via the disk conformal mapping. Figure 1 displays two disk-parameterized surfaces in the middle and right panels; we expand on the differences between these two panels in A.7. Each face, $f$, is thus a map $f:D\rightarrow\mathbb{R}^{3}$ with $D=\{r,\theta|0\leq r\leq 1,0\leq\theta<2\pi\}$, the unit disk.
@@ 3 Novel Face Radial Curves Representation
We parameterize the entire face using a disk, however, our goal is to work with what we refer to as “face radial curves.” This concept is similar to that of facial curves (Samir et al., 2006) and radial curves (Drira et al., 2010). In the facial recognition literature, for instance, to construct a radial curve representation one picks a focal point, typically the tip of the nose, and overlays curves on the surface of the face with the nose as the center and each point of the curve being equidistant from the tip of the nose. That is, given a focal point $p$ each radial curve is $g=\{x|r=d_{F}(p,x)\}$ for a given $r\geq 0$ and where $d_{F}$ is distance measured on the surface of the face.
These previously mentioned methodologies do not take into account that the distribution of features on faces is not uniform for everyone. That is, let $g_{ij}$ represent the $j$th radial curve of the $i$th face, then $g_{ij}$ may be the curve which goes directly on the middle of the eyes for individual $i$ but $g_{i^{\prime}j}$ may be directly on the forehead of individual $i^{\prime}$ for the same $j$. Features, such as eyes and the mouth, being disproportionately farther or closer to the focal point or each other hence causes issues. These methodologies do, however consider registration within each $j$ but not across all $j$s simultaneously. We propose a new method to have an entire global alignment and registration across all sets of curves and faces.
@@
We apply our proposed method to data described in Sero et al. (2019); more details are available in A.4. The left panel of Figure 2 displays a face as a point cloud and the right panel is the same face but with the disk parameterization and some of the face radial curves overlain. While the disk parameterization is an infinite collection of these curves, we pick some number of these curves to represent the face. The number of curves is a tuning parameter where more curves implies, to an extent, more definition in the face. We discuss this further in 5.
@@ 4 Gaussian Differential Privacy for Functional Data
Since the conception of DP (Dwork et al., 2006), noise calibration has been considered from different perspectives leading to many variants. For instance, “zero concentrated differential privacy” (Bun & Steinke, 2016) and “Rényi differential privacy” (Mironov, 2017) are notions of DP from the perspective of a divergence of the distribution of the mechanism. In the present paper we utilize “Gaussian DP” (GDP, Dong et al. (2019)), which considers DP from the perspective of a particular set of hypothesis tests. As one of our key contributions in this paper, we extend GDP to functional data analysis (FDA) in Section 4.2. We begin with a brief overview of DP for FDA.
@@ 4.1 Background on Differential Privacy for Functional Data
In this section and in A.5, we describe approximate DP, $(\epsilon,\delta)$-DP, in the context of FDA. This is a needed background for our proposed extension of GDP into the space of functional data. For a more thorough exposition on DP, see Blum et al. (2005); Dwork & Roth (2014); Dwork et al. (2006).
@@
We consider releasing an estimate of a private mean function. To sanitize functions we rely on tools and foundations of functional data analysis in the domain of privacy as in Hall et al. (2013); Alda & Rubinstein (2017); Mirshani et al. (2017), with the latter considering spaces more extensive than functions. The infinite dimensional nature of function valued summaries presents a critical challenge in establishing formal privacy. In particular, traditional probability densities become much more complicated as there is no default or baseline measure in infinite dimensions (unlike Lebesgue measure in $\mathbb{R}^{d}$). To overcome this challenge, there are currently two approaches. The first, taken by Hall et al. (2013) is to work in finite dimensions and then take careful limits. The second approach, introduced in Mirshani et al. (2017) and which we follow here, is to utilize carefully constructed infinite dimensional densities so that the probability inequalities can be worked out directly (and avoid having to take limits).
@@
Now that we have well defined noise, we consider releasing a specific summary statistic: a private mean function with sample mean $h(D)=\bar{X}=\frac{1}{n}\sum_{i}x_{i}$. To have some control on the smoothness of this estimate, we enforce smoothness by relying on an optimization problem with a penalty. That is, we let $h(D)=\operatorname{argmin}_{m\in\mathcal{H}}\sum_{i}\|x_{i}-m\|^{2}_{\mathbb{H}}+\phi\|m\|_{\mathcal{H}}^{2}$ with $\phi$ being a smoothness penalty parameter. This approach ensures the required noise compatibility. For our summary statistic, the sensitivity, $\sup_{D\sim D^{\prime}}\|h(D)-h(D^{\prime})\|^{2},$ can be shown to be bounded as $\Delta^{2}\leq 4\tau^{2}/(n^{2}\phi)$ (Mirshani et al., 2017), where $\tau$ is a finite bound on the norm in the $\mathbb{H}$ space of the elements of all $D\in\mathcal{D}$.
@@ 4.2 Extension of GDP to Functional Data
One especially useful interpretation of approximate DP comes from Wasserman & Zhou (2010), which relates DP to hypothesis testing. In particular, for a sanitized output, $\tilde{h}(D)$, one can consider statistical tests for determining if the true underlying data source is $D$ or some adjacent data set $D^{\prime}$. In these simple cases, the optimal test is well known (Neyman-Pearson Lemma), so one can talk about the optimal type 2 error (1-power) of a statistical test for given type 1 error rate. It turns out that DP gives a bound for the type 2 error relative to the type 1 error, which Dong et al. (2019) took a step further to define Gaussian DP. In particular, Dong et al. (2019) show a mechanism is $\mu$-GDP if its entire type 1/type 2 error tradeoff curve is bounded by that of a curve coming from comparing a $N(0,1)$ to $N(\mu,0)$, meaning it is harder to distinguish 00 and $\mu$ from $N(0,1)$ than it is $h(D)$ from $h(D^{\prime})$ from $\tilde{h}(D)$. We provide the formal definition below.
@@
Corollary 4.2 also appears in Balle & Wang (2018) in the context of the calibrating the Gaussian mechanism, but in the presented form the corollary is more readily applicable for our setting.
@@
The last equality is a reparametrization setting $\mu=\Delta/\sigma$ and the rest of the details are in the A.8. ∎
For facial radial curves and other disc-like surfaces, by construction, we need to sanitize many mean curves, so we require composition of a multitude of mechanisms with their respective budgets $\mu_{i}$’s. Composition in the GDP framework is “tight.” Given two mechanisms $\tilde{h}_{1}$, $\tilde{h}_{2}$ with privacy parameters $\mu_{1}$ and $\mu_{2}$, their composition $(\tilde{h}_{1},\tilde{h}_{2})$ is $\sqrt{\mu_{1}^{2}+\mu_{2}^{2}}$-GDP (Dong et al., 2019, Corallary 3.3).
@@ 5 Face Radial Curve Example
By design we have closed curves, so to retain this structure we use a periodic kernel that takes the form $k(s,t)=\exp\left(-\left[d_{\mathbb{S}^{1}}(\omega(t),\omega(s))/\rho\right]^{\alpha}\right)$ (Gneiting, 2013) with $t,s\in[0,1]$. The distance of two points $a,b$ on $\mathbb{S}^{1}$ is $\arccos\langle a,b\rangle$, however, the previous $t,s$ are parameters on the unit interval, so we first “wrap” the interval around the circle to compute this distance hence the need for $\omega(\cdot)$, a wrapping function. This kernel is a powered exponential on spheres with kernel range parameter $\rho$ and smoothness parameter $\alpha\in(0,1]$. For our experiments we set $\alpha=1$ as we can control smoothness instead via $\phi$ in the mean estimation as in 4.1.